A rectangle is 4 units wide. How many square units does the rectangle cover if it is 6,7,8, or 9 units long?
Answer by jim_thompson5910(35256)

6 units long: Area = 4*6 = 24 square units


7 units long: Area = 4*7 = 28 square units


I'll let you do the last two.


If you need more help, email me at jim_thompson5910@hotmail.com

Also, feel free to check out my tutoring website

Jim